# Unlearning as multi-task optimization: A normalized gradient difference approach with an adaptive learning rate

### v1

@@ Abstract
[…]
[table]capposition=above
Unlearning as multi-task optimization:
A normalized gradient difference approach with an adaptive learning rate
Zhiqi Buthanks: Equal contribution. Work done during Xiaomeng Jin’s internship at Amazon. Amazon zhiqibu@amazon.com                        Xiaomeng Jin UIUC                        Bhanukiran Vinzamuri Amazon                        Anil Ramakrishna Amazon
Kai-Wei Changthanks: Concurrent positions as an Amazon Scholar and as a faculty at the corresponding institutes. UCLA                        Volkan Cevher EPFL                        Mingyi Hong University of Minnesota
@@ 1 Introduction
Existing machine unlearning methods are formulated primarily as minimizing memorization through the language model loss Jang et al. (2023); Chen et al. (2024); Liu et al. (2024b). In particular, the Gradient Ascent (GA) method maximizes the language model (LM) loss (i.e., minimizes the negative LM loss) on the target forget set (F). However, this approach can also negatively affect the utility of the model. To mitigate the utility loss, the Gradient Difference (GDiff) method selects a subset of the training data as the retain set (R), minimizing the sum of the negative LM loss on the forgetting set and the standard LM loss on the retaining set. This approach has been empirically shown to effectively preserve the model’s performance Liu et al. (2022); Maini et al. (2024). Similarly, Negative Preference Optimization (NPO) Zhang et al. (2024b) assigns a lower likelihood of forgetting data, thereby balancing the unlearning performance with model utility.
@@
Despite these successes, there are still two key issues preventing the methods from reaching their full potential. First, balancing retaining and forgetting losses is difficult (Figure 1, details are in Appendix A) given the disproportionate sizes of the forget and retain datasets. Second, the optimization methods for unlearning are usually sensitive to the learning rate (cf., Appendix A, Figure 7). For instance, various learning rates can lead to substantial changes in the ROUGE scores and loss values even for the same algorithm, making the unlearning methods unstable and difficult to use in practice.
@@
In this paper, we carefully examine unlearning from an optimization perspective and formulate it as a multi-task optimization (MTO) problem Chen et al. (2021); Xin et al. (2022): we aim to minimize the LM loss (i.e., maximize the utility) on the retaining set and maximize the LM loss on the forgetting set (i.e., minimize memorization), simultaneously. [^1] A naive approach is optimizing the sum of these two objectives. We will discuss alternatives to improve upon this. To solve this two-task problem, we study the rich literature of multi-task methods that seeks the Pareto optimality of two tasks (e.g., IMTL Liu et al. (2021), GradNorm Chen et al. (2018a), RLW Lin et al. (2021), PCGrad Yu et al. (2020), and scalarization Boyd and Vandenberghe (2004)), and design an approach specifically for the LLM unlearning problem.
Figure 1: Memorization (measured by ROUGE) on TOFU forgetting and retaining sets with Phi-1.5 (see setup in Sec 5). The dashed line represents extended GDiff with $0\leq c\leq 1$ in (3). Our NGDiff (gray square) achieves the best trade-off (bottom right preferred) among other unlearning methods.
@@
Inspired by the simplicity and strong empirical performance of linear scalarization methods, [^2] Xin et al. (2022) demonstrate that linear scalarization outperforms, or is at least on par with, other MTO approaches across various language and vision experiments which minimize a linearly weighted average of task losses, we propose an LLM unlearning method, NGDiff, based on dynamic scalarization, and analyze its theoretical properties. Building on the analysis, we introduce an automatic learning rate adaptation method tailored for LLM unlearning.
@@
We showcase the effectiveness of our method through extensive experiments on multiple datasets, different LLMs and vision models. For example, on TOFU Maini et al. (2024), NGDiff achieves 40% higher model utility while maintaining comparable unlearning performance with Llama2-7B. Figure 1 highlights the effectiveness of NGDiff.
@@
  • We integrate NGDiff with GeN Bu and Xu (2024), which uses Hessian-based learning rate selection for stable convergence.
@@ 2 Related Work
LLM unlearning The extensive data used in training LLMs raises significant concerns. Certain data sources contain personal information Carlini et al. (2021), outdated knowledge Wu et al. (2024), and copyright-protected materials Times (2023). In addition, adversarial data attacks can maliciously manipulate training data to embed harmful information Wallace et al. (2021); Li et al. (2024).
To remove unwanted information without retraining the entire model, machine unlearning has been proposed using techniques such as data slicing Bourtoule et al. (2021), influence functions Ullah et al. (2021), and differential privacy Gupta et al. (2021). However, these methods are challenging to scale to LLMs due to their complexity. Recently, efficient approximate unlearning methods have been proposed for LLMs Eldan and Russinovich (2023); Zhang et al. (2024a); Jang et al. (2023); Pawelczyk et al. (2024); Chen and Yang (2023). They mostly focus on designing unlearning objectives or hiding unwanted information. However, none addresses the fundamental optimization problem. Our paper bridges this gap and complements existing approaches. Further discussion of the challenges surrounding LLM unlearning can be found in benchmarks Shi et al. (2024); Maini et al. (2024) and surveys Si et al. (2023); Liu et al. (2024c).
Note that the literature on knowledge editing De Cao et al. (2021) is also relevant. However, model editing typically focuses on surgically updating LLMs for specific knowledge, whereas unlearning removes the influence of particular documents. The techniques presented in this paper could potentially be applied to knowledge editing.
Multi-task optimization In NLP, multi-task learning Zhang et al. (2023) typically refers to building a model that can perform well on multiple tasks simultaneously by sharing representations, introducing constraints, or combining multiple learning objectives [^3] This often involves optimizing a form of static linear scalarization, as introduced in the next section. Multi-task optimization, on the other hand, focuses on a slightly different concept – optimizing two distinct learning objectives simultaneously. The key challenge is how to balance the trade-off among objectives during the optimization procedure by modifying the per-task gradients (e.g. PCGrad Yu et al. (2020), RLW Lin et al. (2021), IMTL Liu et al. (2021)). Several recent works have studied Pareto frontier and optimally in context of NLP tasks (e.g., multi-lingual machine translation Chen et al. (2023); Xin et al. (2022) and NLP fairness Han et al. (2023)). While optimization is often treated as a black-box tool in NLP research, studying optimization provides deeper insights and inspires new algorithms. [^4] An example is the Baum-Welch algorithm, originally proposed to estimate the parameters of HMMs and applied in speech recognition in 1970s before it was recognized as an instance of the EM algorithm Dempster et al. (1977). It was later identified as a special case of a broader class of convex-concave optimization (CCCP) Yuille and Rangarajan (2001). This connection inspired new designs, such as the unified EM Samdani et al. (2012).
@@ 3.1 Static linear scalarization
$\bm{g}_{\text{static}}(\bm{\theta}_{t};\!c)\!\!=\!\!\frac{\partial\textup{LSP}}{\partial\bm{\theta}_{t}}\!\!=\!\!c\bm{g}_{\textup{R}}(\bm{\theta}_{t})\!-\!(1\!\!-\!\!c)\bm{g}_{\textup{F}}(\bm{\theta}_{t}).$ (3)
@@
$\bm{\theta}_{t+1}=\bm{\theta}_{t}-\eta_{t}[c\bm{g}_{\textup{R}}(\bm{\theta}_{t})-(1-c)\bm{g}_{\textup{F}}(\bm{\theta}_{t})].$
@@
We term the static linear scalarization as the extended GDiff in this work. Some existing methods are special cases of extended GDiff. For example, Gradient Descent (GD) on retaining set is equivalent to extended GDiff with $c=1$. Gradient Ascent (GA) on forgetting is equivalent to that with $c=0$, and vanilla GDiff Liu et al. (2022) set $c=0.5$ (i.e., equally weighted).
Figure 2: Gradient space in 2-dimension. $\color[rgb]{1,0,0}\bm{g}_{\textup{F}}$ is the forgetting gradient and $\color[rgb]{0,0,1}\bm{g}_{\textup{R}}$ is the retaining gradient, each with a perpendicular dashed line. Yellow area is the linear span (Eq. (3)) by scalarization. Green area is positively correlated to $\bm{g}_{\textup{R}}$ and negatively correlated to $\bm{g}_{\textup{F}}$ by Eq. (6), whereas NGDiff always stays within at each iteration.
A nice property of linear scalarization is the Pareto optimality at the convergence of convex models Boyd and Vandenberghe (2004)), which we state in Lemma 2 (proof in Appendix G) for the static $c$ and later extend to Theorem 3 for the dynamic $c_{t}$ in Section 3.2.
Lemma 2 (restated from Xin et al. (2022)).
@@ 3.2 Dynamic scalarization
$\bm{g}_{\text{UN}}(\bm{\theta};c_{t})$ $:=c_{t}\bm{g}_{\textup{R}}(\bm{\theta})-(1-c_{t})\bm{g}_{\textup{F}}(\bm{\theta}).$ (4)
@@
    RLW Lin et al. (2021), $c_{t}=\frac{e^{\lambda_{1}}}{e^{\lambda_{1}}+e^{\lambda_{2}}}$ with $\lambda_{i}\sim N(0,1)$
@@
    PCGrad Yu et al. (2020), $\frac{c_{t}}{1-c_{t}}=1+\frac{\bm{g}_{\textup{F}}^{\top}\bm{g}_{\textup{R}}}{|\bm{g}_{\textup{R}}|^{2}}$
@@
    IMTL-G Liu et al. (2021), $c_{t}=\bm{g}_{\textup{F}}^{\top}(\frac{\bm{g}_{\textup{F}}}{||\bm{g}_{\textup{F}}||}-\frac{\bm{g}_{\textup{R}}}{||\bm{g}_{\textup{R}}||})/(\bm{g}_{\textup{F}}-\bm{g}_{\textup{R}})^{\top}(\frac{\bm{g}_{\textup{F}}}{||\bm{g}_{\textup{F}}||}-\frac{\bm{g}_{\textup{R}}}{||\bm{g}_{\textup{R}}||})$
@@ 4.1 Loss landscape of unlearning
Applying the Taylor expansion on Eq. (4), we can view the local landscapes of loss $L_{\textup{R}}$ and $L_{\textup{F}}$ as quadratic functions, where $L_{\omega}(\bm{\theta}_{t+1})-L_{\omega}(\bm{\theta}_{t})=$
$-\eta_{t}\bm{g}_{\omega}^{\top}\bm{g}_{\text{UN}}(c_{t})+(\eta_{t}^{2}/2)\bm{g}_{\text{UN}}^{\top}\mathbf{H}_{\omega}\bm{g}_{\text{UN}}+o(\eta_{t}^{2}),$ (5)
@@
Here $\mathbf{H}_{\omega}=\frac{\partial^{2}L_{\omega}}{\partial\bm{\theta}^{2}}$ is the Hessian matrix, which empirically gives $\bm{g}_{\text{UN}}^{\top}\mathbf{H}_{\omega}\bm{g}_{\text{UN}}>0$ and renders $L_{\textup{R}}$ and $L_{\textup{F}}$ locally and directionally convex along the gradients. This allows the existence of a minimizing learning rate to be characterized in Section 4.3. We visualize the loss landscape of Phi-1.5 Li et al. (2023a) model on an unlearning benchmark, TOFU dataset Maini et al. (2024) in Figure 3 and observe that the quadratic functions in Eq. (5) are well-fitted in most iterations.
@@ 4.2 Normalized gradient difference
To satisfy Eq. (6) , we propose a normalized gradient difference method (NGDiff) to dynamically set $c_{t}=\frac{1/||\bm{g}_{\textup{R}}||}{1/||\bm{g}_{\textup{R}}||+1/||\bm{g}_{\textup{F}}||}\Longrightarrow\bm{g}_{\textup{NGDiff}}(\bm{g}_{\textup{R}},\bm{g}_{\textup{F}}):=\frac{\bm{g}_{\textup{R}}}{||\bm{g}_{\textup{R}}||}-\frac{\bm{g}_{\textup{F}}}{||\bm{g}_{\textup{F}}||}.$ In words, we normalize the retaining and forgetting gradients [^6] We illustrate in Appendix B that NGDiff is critically different and simpler than GradNorm.(Chen et al., 2018a).
@@
The condition, $\bm{g}_{\textup{NGDiff}}^{\top}\mathbf{H}\bm{g}_{\textup{NGDiff}}>0$ in part (2) of Theorem 5, may not always hold in deep learning. However, it empirically holds in most iterations across models and datasets in our experiments (cf., our Figure 3 and Figure 2 in Bu and Xu (2024)), and we can stablize the training by not updating $\eta$ when the condition fails.
@@ 4.3 Automatic learning rate adaption
In order for NGDiff to work as in Theorem 5, the learning rate schedule needs to be carefully selected so that $0<\eta_{t}<\frac{2\bm{g}_{\textup{R}}^{\top}\bm{g}_{\textup{NGDiff}}}{\bm{g}_{\textup{NGDiff}}^{\top}\mathbf{H}_{\textup{R}}\bm{g}_{\textup{NGDiff}}}$ at each iteration. In Algorithm 1, we adapt GeN (Bu and Xu, 2024) (or AutoLR) to the unlearning setting and dynamically set the learning rates [^7] We note other parameter-free methods such as D-adaptation, Prodigy, and DoG can also set the learning rate automatically. However, these methods need to be tailored for different gradient methods, hence not compatible to NGDiff or the unlearning algorithms in general. We give a detailed explanation in Appendix E. as the minimizer of (5): to locally optimize $L_{\textup{R}}$ and to monotonically increase $L_{\textup{F}}$, we use the following learning rate:
@@
Algorithm 1 summarizes NGDiff.
@@
7:     Construct $\bm{g}_{\textup{NGDiff}}=\bm{g}_{\textup{R}}/||\bm{g}_{\textup{R}}||-\bm{g}_{\textup{F}}/||\bm{g}_{\textup{F}}||$
@@
11:         Fit the quadratic function in Eq. (5) from $(-\eta,0,\eta)\to(L_{\textup{R}}^{-},L_{\textup{R}},L_{\textup{R}}^{+})$
12:         Derive the optimal learning rate $\eta_{t}^{*}$ by Eq. (7) and set $\eta=\eta_{t}^{*}$

### v2

@@ 1 Introduction
Existing machine unlearning methods are formulated primarily as minimizing memorization through the language model loss [22, 7, 34]. In particular, the Gradient Ascent (GA) method maximizes the language model (LM) loss (i.e., minimizes the negative LM loss) on the target forget set (F). However, this approach can also negatively affect the utility of the model. To mitigate the utility loss, the Gradient Difference (GDiff) method selects a subset of the training data as the retain set (R), minimizing the sum of the negative LM loss on the forgetting set and the standard LM loss on the retaining set. This approach has been empirically shown to effectively preserve the model’s performance [31, 36]. Similarly, Negative Preference Optimization (NPO) [55] assigns a lower likelihood of forgetting data, thereby balancing the unlearning performance with model utility.
@@
Despite these successes, there are still two key issues preventing the methods from reaching their full potential. First, balancing retaining and forgetting losses is difficult (Figure 1) given the disproportionate sizes of the forget and retain datasets. Second, the optimization methods for unlearning are usually sensitive to the learning rate (cf., Appendix A, Figure 6). For instance, various learning rates can lead to substantial changes in the ROUGE scores and loss values even for the same algorithm, making the unlearning methods unstable and difficult to use in practice.
@@
In this paper, we carefully examine unlearning from an optimization perspective and formulate it as a multi-task optimization (MTO) problem [9, 51]: we aim to minimize the LM loss (i.e., maximize the utility) on the retaining set and maximize the LM loss on the forgetting set (i.e., minimize memorization), simultaneously. [^1] A naive approach is optimizing the sum of these two objectives. We will discuss alternatives to improve upon this. To solve this two-task problem, we study the rich literature of multi-task methods that seeks the Pareto optimality of two tasks (e.g., IMTL [33], GradNorm [10], RLW [30], PCGrad [52], and scalarization [3]), and design an approach specifically for the LLM unlearning problem.
Figure 1: Loss values and ROUGE scores on the forgetting and retaining data from the TOFU dataset using different unlearning methods on the Phi-1.5 language model. We apply the extended GDiff with various coefficients (see (3), $0\leq c\leq 1$) and connect the results with a blue dashed line. We denote MTO methods as different markers, and use a grey dashed line to represent the loss of random guess.
@@
Inspired by the simplicity and strong empirical performance of linear scalarization methods,222xin2022current [51] demonstrate that linear scalarization outperforms, or is at least on par with, other MTO approaches across various language and vision experiments which minimize a linearly weighted average of task losses, we propose an LLM unlearning method, NGDiff, based on dynamic scalarization, and analyze its theoretical properties. Building on the analysis, we introduce an automatic learning rate adaptation method tailored for LLM unlearning.
@@
We showcase the effectiveness of our method through extensive experiments on multiple datasets, different LLMs and vision models. For example, on TOFU [36], NGDiff achieves 40% higher model utility while maintaining comparable unlearning performance with Llama2-7B. Figure 1 highlights the effectiveness of NGDiff.
@@
  • We integrate NGDiff with GeN [4], which uses Hessian-based learning rate selection for stable convergence.
@@ 2 Related Work
LLM unlearning The extensive data used in training LLMs raises significant concerns. Certain data sources contain personal information [5], outdated knowledge [50], and copyright-protected materials [45]. In addition, adversarial data attacks can maliciously manipulate training data to embed harmful information [48, 27].
To remove unwanted information without retraining the entire model, machine unlearning has been proposed using techniques such as data slicing [2], influence functions [47], and differential privacy [18]. However, these methods are challenging to scale to LLMs due to their complexity. Recently, efficient approximate unlearning methods have been proposed for LLMs [17, 54, 22, 39, 6]. They mostly focus on designing unlearning objectives or hiding unwanted information. However, none addresses the fundamental optimization problem. Our paper bridges this gap and complements existing approaches. Further discussion of the challenges surrounding LLM unlearning can be found in benchmarks [43, 36] and surveys [44, 35].
Note that the literature on knowledge editing [13] is also relevant. However, model editing typically focuses on surgically updating LLMs for specific knowledge, whereas unlearning removes the influence of particular documents. The techniques presented in this paper could potentially be applied to knowledge editing.
Multi-task optimization In NLP, multi-task learning [56] typically refers to building a model that can perform well on multiple tasks simultaneously by sharing representations, introducing constraints, or combining multiple learning objectives [^3] This often involves optimizing a form of static linear scalarization, as introduced in the next section. Multi-task optimization, on the other hand, focuses on a slightly different concept – optimizing two distinct learning objectives simultaneously. The key challenge is how to balance the trade-off among objectives during the optimization procedure by modifying the per-task gradients (e.g. PCGrad [52], RLW [30], IMTL [33]). Several recent works have studied Pareto frontier and optimally in context of NLP tasks (e.g., multi-lingual machine translation [8, 51] and NLP fairness [19]). While optimization is often treated as a black-box tool in NLP research, studying optimization provides deeper insights and inspires new algorithms. [^4] An example is the Baum-Welch algorithm, originally proposed to estimate the parameters of HMMs and applied in speech recognition in 1970s before it was recognized as an instance of the EM algorithm [15]. It was later identified as a special case of a broader class of convex-concave optimization (CCCP) [53]. This connection inspired new designs, such as the unified EM [42].
@@ 3.1 Static linear scalarization
$\bm{g}_{\text{static}}(\bm{\theta}_{t};c)=\frac{\partial\textup{LSP}}{\partial\bm{\theta}_{t}}=c\cdot\bm{g}_{\textup{R}}(\bm{\theta}_{t})-(1-c)\cdot\bm{g}_{\textup{F}}(\bm{\theta}_{t}).$ (3)
@@
$\bm{\theta}_{t+1}=\bm{\theta}_{t}-\eta_{t}[c\cdot\bm{g}_{\textup{R}}(\bm{\theta}_{t})-(1-c)\cdot\bm{g}_{\textup{F}}(\bm{\theta}_{t})].$ (4)
@@
We term the static linear scalarization as the extended GDiff in this work. Some existing methods are special cases of extended GDiff. For example, Gradient Descent (GD) on retaining set is equivalent to extended GDiff with $c=1$. Gradient Ascent (GA) on forgetting is equivalent to that with $c=0$, and vanilla GDiff [31] set $c=0.5$ (i.e., equally weighted).
Figure 2: Gradient space in 2-dimension. $\color[rgb]{1,0,0}\bm{g}_{\textup{F}}$ is the forgetting gradient and $\color[rgb]{0,0,1}\bm{g}_{\textup{R}}$ is the retaining gradient, each with a perpendicular dashed line. Yellow area is the linear span (Eq. (3)) by scalarization. Green area is positively correlated to $\bm{g}_{\textup{R}}$ and negatively correlated to $\bm{g}_{\textup{F}}$ by Eq. (7), whereas NGDiff always stays within this green area at each iteration.
A nice property of linear scalarization is the Pareto optimality at the convergence of models, which we state in Lemma 2 (proof in Appendix G) for the static $c$ and later extend to Theorem 3 for the dynamic $c_{t}$ in Section 3.2.
Lemma 2 (restated from [51]).
For any $10<c<$ 1, the model $\bm{\theta}^{*}_{\text{LSP}}(c)\in\operatorname*{arg\,min}_{\theta}\textup{LSP}(\bm{\theta};c)$ is Pareto optimal.
@@ 3.2 Dynamic scalarization
$\bm{g}_{\text{UN}}(\bm{\theta};c_{t})$ $:=c_{t}\cdot\bm{g}_{\textup{R}}(\bm{\theta})-(1-c_{t})\cdot\bm{g}_{\textup{F}}(\bm{\theta}).$ (5)
@@
    RLW [30], $c_{t}=\frac{e^{\lambda_{1}}}{e^{\lambda_{1}}+e^{\lambda_{2}}}$ with $\lambda_{i}\sim N(0,1)$
@@
    PCGrad [52], $\frac{c_{t}}{1-c_{t}}=1+\frac{\bm{g}_{\textup{F}}^{\top}\bm{g}_{\textup{R}}}{\|g_{\textup{R}}\|^{2}}$
@@
    IMTL-G [33], $c_{t}=\bm{g}_{\textup{F}}^{\top}(\frac{\bm{g}_{\textup{F}}}{\|\bm{g}_{\textup{F}}\|}-\frac{\bm{g}_{\textup{R}}}{\|\bm{g}_{\textup{R}}\|})/(\bm{g}_{\textup{F}}-\bm{g}_{\textup{R}})^{\top}(\frac{\bm{g}_{\textup{F}}}{\|\bm{g}_{\textup{F}}\|}-\frac{\bm{g}_{\textup{R}}}{\|\bm{g}_{\textup{R}}\|})$
@@
For any $\{c_{t}\}$ with $0\leq c_{t}\leq 1$ that converges as $t\to\infty$, the model $\bm{\theta}^{*}(\{c_{t}\}):=\lim_{t\to\infty}\bm{\theta}_{t}$ in (5) is Pareto optimal.
@@ 4.1 Loss landscape of unlearning
Applying the Taylor expansion on Eq. (5), we can view the local landscapes of loss $L_{\textup{R}}$ and $L_{\textup{F}}$ as quadratic functions, where
$L_{\omega}(\bm{\theta}_{t+1})-L_{\omega}(\bm{\theta}_{t})=-\eta_{t}\bm{g}_{\omega}^{\top}\bm{g}_{\text{UN}}(c_{t})+(\eta_{t}^{2}/2)\bm{g}_{\text{UN}}^{\top}\mathbf{H}_{\omega}\bm{g}_{\text{UN}}+o(\eta_{t}^{2}),$ (6)
@@
Here $\mathbf{H}_{\omega}=\frac{\partial^{2}L_{\omega}}{\partial\bm{\theta}^{2}}$ is the Hessian matrix, which empirically gives $\bm{g}_{\text{UN}}^{\top}\mathbf{H}_{\omega}\bm{g}_{\text{UN}}>0$ and renders $L_{\textup{R}}$ and $L_{\textup{F}}$ locally and directionally convex along the gradients. This allows the existence of a minimizing learning rate to be characterized in Section 4.3. We visualize the loss landscape of Phi-1.5 [28] model on an unlearning benchmark, TOFU dataset [36] in Figure 3 and observe that the quadratic functions in Eq. (6) are well-fitted in most iterations.
@@ 4.2 Normalized gradient difference
To satisfy Eq. (7) , we propose a normalized gradient difference method (NGDiff) to dynamically set $c_{t}=\frac{1/\|\bm{g}_{\textup{R}}\|}{1/\|\bm{g}_{\textup{R}}\|+1/\|\bm{g}_{\textup{F}}\|}\Longrightarrow\bm{g}_{\textup{NGDiff}}(\bm{g}_{\textup{R}},\bm{g}_{\textup{F}}):=\frac{\bm{g}_{\textup{R}}}{\|\bm{g}_{\textup{R}}\|}-\frac{\bm{g}_{\textup{F}}}{\|\bm{g}_{\textup{F}}\|}.$ In words, we normalize the retaining and forgetting gradients [^6] We illustrate in Appendix B that NGDiff is critically different and simpler than GradNorm.[10].
@@
The condition, $\bm{g}_{\textup{NGDiff}}^{\top}\mathbf{H}\bm{g}_{\textup{NGDiff}}>0$ in part (2) of Theorem 5, may not always hold in deep learning. However, it empirically holds in most iterations across models and datasets in our experiments (cf., our Figure 3 and Figure 2 in [4]), and we can stablize the training by not updating $\eta$ when the condition fails.
@@ 4.3 Automatic learning rate adaption
In order for NGDiff to work as in Theorem 5, the learning rate schedule needs to be carefully selected so that $0<\eta_{t}<\frac{2\bm{g}_{\textup{R}}^{\top}\bm{g}_{\textup{NGDiff}}}{\bm{g}_{\textup{NGDiff}}^{\top}\mathbf{H}_{\textup{R}}\bm{g}_{\textup{NGDiff}}}$ at each iteration. In Algorithm 1, we adapt GeN [4] (or AutoLR) to the unlearning setting and dynamically set the learning rates [^7] We note other parameter-free methods such as D-adaptation, Prodigy, and DoG can also set the learning rate automatically. However, these methods need to be tailored for different gradient methods, hence not compatible to NGDiff or the unlearning algorithms in general. We give a detailed explanation in Appendix E. as the minimizer of (6): to locally optimize $L_{\textup{R}}$ and to monotonically increase $L_{\textup{F}}$, we use the following learning rate:
@@
Algorithm 1 summarizes NGDiff. We note that $\bm{g}_{\textup{F}},\bm{g}_{\textup{R}}$ can be stochastic gradients.
@@
7:     Construct $\bm{g}_{\textup{NGDiff}}=\bm{g}_{\textup{R}}/\|\bm{g}_{\textup{R}}\|-\bm{g}_{\textup{F}}/\|\bm{g}_{\textup{F}}\|$
@@
11:         Fit the quadratic function $\mathcal{Q}$ in Eq. (6) from $(-\eta,0,\eta)\to(L_{\textup{R}}^{-},L_{\textup{R}},L_{\textup{R}}^{+})$
12:         Derive the optimal learning rate $\eta_{t}^{*}:=\text{argmin}_{x}\mathcal{Q}(x)$ and set $\eta=\eta_{t}^{*}$